# From Text to Talk: Audio-Language Model Needs Non-Autoregressive Joint Training

### v1

@@ Abstract
Recent advances in large language models (LLMs) have attracted significant interest in extending their capabilities to multimodal scenarios, particularly for speech-in/speech-out conversational systems. However, existing multimodal models handling interleaved audio and text—such as MOSHI, GLM4-Voice, and VITA-Audio require complex multi-stage training pipelines, incurring substantial computational costs. Moreover, these models uniformly apply autoregressive (AR) generation to both text and audio tokens, overlooking a fundamental asymmetry in their dependency structures: while text tokens exhibit strong target-target dependencies requiring causal ordering, audio tokens are predominantly driven by source-target dependencies, where audio outputs primarily condition on source text rather than preceding audio tokens. In this work, we propose Text-to-Talk (TtT), a unified audio-text modeling framework that integrates AR text generation with non-autoregressive (NAR) audio diffusion within a single Transformer architecture initialized from a pretrained LLM. Leveraging the any-order autoregressive nature of absorbing discrete diffusion, we establish a unified theoretical framework that combines conventional AR for text with NAR diffusion for audio, proving that our joint training objective provides an upper bound on the negative log-likelihood of the desired joint distribution. To support this hybrid generation paradigm, we design a modality-aware attention mechanism that enforces causal decoding for text while enabling bidirectional context modeling within audio spans. To address the inherent train-test discrepancies in hybrid AR-NAR training, we introduce three principled strategies to bridge the gap between training and inference conditions and enable robust content-aware prediction of end-of-audio tokens for variable-length generation. During inference, TtT employs a block-wise diffusion algorithm that enables parallel audio synthesis while maintaining the ability to generate outputs of arbitrary length. Extensive experiments across ASR and Audio-QA tasks demonstrate the effectiveness of our approach, with detailed ablation studies validating each proposed component. We will open-source our models, data, and code to facilitate future research in this direction.
@@ 1 Introduction
The recent success of LLMs has catalyzed a paradigm shift towards general-purpose Multimodal Large Language Models (MLLMs) capable of processing and generating information across diverse modalities (Xu et al., 2025; Team et al., 2023). Among these, speech-in/speech-out conversational systems have emerged as a pivotal component in facilitating natural human-AI interaction. Conventional systems typically decompose this problem into a cascaded pipeline of automatic speech recognition (ASR), text-based language model inference, and text-to-speech (TTS) synthesis. While effective to a degree, this modular design introduces significant latency accumulation and critical error propagation between modules, hindering naturalness and real-world applicability. In response, recent end-to-end approaches like Moshi Défossez et al. (2024), GLM4-Voice Zeng et al. (2024), and VITA-Audio Long et al. (2025) have sought to unify speech understanding and generation within a single model. As shown in Fig. 1, these methods aim to generate interleaved text and speech tokens in an autoregressive manner, which are then decoded into continuous audio waveforms by a separate neural codec or diffusion-based decoder (Mehta et al., 2024; Kong et al., 2020).
Figure 1: Challenges in conventional AR audio-text MLLMs. (a) AR models apply uniform autoregressive objectives to both text and audio, overlooking their distinct dependency structures: text exhibits target-target dependencies while audio primarily depends on source-target relationships. (b) Due to disparate tokenization rates, text tokens are exhausted early in the sequence. The preceding audio spans are kept at a consistent size, whereas the final audio span absorbs the remainder and is of variable length, which introduces a strong positional bias, complicating end-of-span prediction for NAR decoders.
@@
However, this emerging paradigm faces two significant challenges. First, existing audio-text MLLMs require complex multi-stage training pipelines involving text-to-audio tokenizer training, interleaved data construction, text-audio alignment and task-oriented supervised fine-tuning (Huang et al., 2025; Li et al., 2025; Ding et al., 2025; Chu et al., 2024), incurring substantial computational costs. Second, as illustrated in Figure 1, we identify a fundamental mismatch in prevailing approaches that employ a single language model to autoregressively generate both text and audio tokens (Zeng et al., 2024; Xie & Wu, 2024; Borsos et al., 2023; Dang et al., 2024; Rubenstein et al., 2023). This uniform treatment applies identical autoregressive training objectives across both modalities, overlooks a critical distinction in their underlying generative processes. Text generation inherently follows a sequential causal structure characterized by strong target-target dependencies Box et al. (2015), where each token explicitly conditions on previously generated tokens. Consequently, an incorrect token prediction can propagate and introduce subsequent errors due to the exposure bias inherent in AR models Ranzato et al. (2015). In contrast, audio token generation is predominantly driven by source-target dependencies Ren et al. (2020), where audio output primarily condition on the source text rather than on the preceding audio tokens. Specifically, within the current NAR span, audio tokens generation should remain faithful to the source text even when previous audio tokens are incorrectly predicted. Applying a purely AR objective to audio generation thus introduces unnecessary sequential constraints, leading to suboptimal training dynamics and magnifying error propagation. This problem can be substantially alleviated by adopting a non-autoregressive generation strategy, which aligns better with the source-dependent nature of audio modeling.
Recently, discrete diffusion has emerged as a compelling alternative to AR for discrete sequence modeling (Yu et al., 2025; Gong et al., 2024; Austin et al., 2021; Sahoo et al., 2024). Beyond empirical gains, recent theory shows that absorbing discrete diffusion can be interpreted as modeling the conditional distributions of clean tokens and admits a tight connection to any‑order AR objectives Ou et al. (2024). Complementarily, information‑theoretic analyses provide convergence guarantees for diffusion language models—sampling error decreases as $\mathcal{O}(1/T)$ with steps and scales with mutual information Li & Cai (2025). Practically, adapting pretrained AR LMs into diffusion LMs offers a scalable and data‑efficient path compared to training diffusion from scratch (Gong et al., 2024). These insights motivate revisiting audio generation with NAR diffusion while preserving AR text inside a single backbone.
Thus, we introduce Text‑to‑Talk (TtT), a unified audio‑text MLLM that integrates AR text generation with NAR audio diffusion within a single Transformer initialized from a pretrained LLM. Leveraging the any-order autoregressive nature of absorbing discrete diffusion, we establish a unified theoretical framework that combines conventional AR for text with NAR diffusion for audio, proving that our joint training objective provides an upper bound on the negative log-likelihood of the desired joint distribution. This design enables the model to autoregressively produce textual tokens while generating corresponding audio spans in parallel via discrete diffusion, subsequently decoded to waveforms with an audio decoder. During inference, TtT employs a block‑by‑block NAR denoising schedule, enabling flexible‑length audio token generation and improved efficiency.
The hybrid nature of our framework introduces a significant train-test discrepancy. To address this, we propose three principled training strategies: (1) Batchwise AR–NAR Objective Mixing (BANOM), which ensures text spans can observe unmasked audio prefixes during training for a portion of samples; (2) Prefix Preservation Masking (PPM), which allows later audio spans to condition on clean preceding audio spans for a fraction of training samples; and (3) Stochastic Span Truncation (SST), which enables content-aware end-of-audio prediction by randomly truncating audio spans during training. These strategies collectively bridge the distributional gap between training and inference, enabling robust variable-length generation.
In summary, our work makes the following contributions:
@@
  • We propose TtT, a hybrid AR-NAR MLLM that seamlessly integrates autoregressive text generation with discrete diffusion-based audio synthesis within a single Transformer initialized from a pretrained LLM. Our design preserves the reasoning and instruction-following capabilities of the base LLM while enabling efficient parallel audio generation through single-stage training.
@@ 2 Preliminary and notation
In this section, we establish the basic notation for interleaved audio-text sequences and provide brief overviews of the two core generative paradigms employed in our framework: AR modeling and absorbing discrete diffusion. These serve as the theoretical foundation upon which our proposed method in Section 3 is built.
@@ 3.3 Multimodal Factorization and a Unified Objective
Intuitively, this averages the likelihood over all possible within-span orderings, reflecting the fact that audio tokens can be generated in any order given the appropriate cross-modal context. Note that while tokens within $\mathcal{A}_{m}$ form an antichain under the partial order (i.e., no mandatory sequential constraints), the order-marginalized conditional in Eq. 13 allows the model to leverage local target-target dependencies that may arise under specific generation orders. This flexibility enables the model to capture useful intra-span structures when beneficial, while preserving the ability to generate in parallel when such dependencies are weak or absent.
@@
This formulation makes explicit that both modalities are fundamentally autoregressive: text uses a single linear extension (left-to-right), while audio integrates over all linear extensions consistent with the partial order.
@@
The key theoretical insight is that our combined training objective provides a tight upper bound on the negative log-likelihood of the desired joint distribution. To see this, consider the audio term. By Jensen’s inequality:
@@
As illustrated in Figure 2 (a), our training pipeline starts from a pretrained text LLM and expands its vocabulary with discrete audio codebook tokens and control symbols ($\langle\text{SOA}\rangle$, $\langle\text{EOA}\rangle$). Each training sequence is organized as interleaved text spans and audio spans, where text spans are optimized with AR Loss while audio spans use the NAR diffusion objective. The unified loss function combines both objectives:
$\mathcal{L}_{\text{Unified}}(x)=\sum_{m=1}^{M}\left[\mathcal{L}_{\text{AR}}(\mathcal{T}_{m})+\beta\cdot\mathcal{L}_{\text{NAR}}(\mathcal{A}_{m})\right]$ (21)
where $\beta$ balances the gradient magnitudes between modalities. However, the hybrid AR-NAR paradigm introduces significant train-test discrepancies that can degrade generation quality. During training, audio spans are partially masked according to the diffusion process, while during inference, the model must generate audio and text tokens conditioned on complete text context and previously generated clean audio tokens. To bridge this distributional gap, we propose three principled training strategies:
@@ 4.2 Main Results
Table 1 main results part reports the comparison of conversational understanding ability across different models on Audio-QA task. From these results we make the following observations: (1) Our proposed TtT consistently achieves the best performance on all four Audio-QA datasets. For example, at the 3B scale, TtT attains scores of 17.46, 34.68, 6.53, and 11.61 on AlpacaEval, LLaMAQuestions, TriviaQA, and WebQuestions respectively, surpassing the Qwen2.5-Base (AR) model by +3.04, +24.68, +5.93, and +10.91. This improvement stems from the hybrid AR–NAR design, which better captures the inherent asymmetry between text and audio dependencies. (2) Our results demonstrate a clear scaling trend shared with the baselines, where larger backbones consistently yield stronger performance. For instance, TtT-3B achieves 34.68 on LLaMAQuestions, outperforming its 1.5B counterpart which scores 23.75, reflecting an absolute gain of +10.93. This pattern aligns with the scaling laws of LLMs, reflecting that additional capacity enhances both reasoning ability and cross-modal representation learning; (3) Qwen2.5-Base (NAR) performs noticeably worse than both our hybrid framework and the AR-only baseline. The performance gap highlights the limitations of purely NAR generation for conversational tasks that require long-range coherence and reasoning. Without AR text modeling, the NAR backbone struggles to maintain causal dependencies across spans, leading to degraded question answering performance.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have attracted significant interest in extending their capabilities to multimodal scenarios, particularly for speech-to-speech conversational systems. However, existing multimodal models handling interleaved audio and text rely on autoregressive methods, overlooking that text depends on target-target relations whereas audio depends mainly on source-target relations. In this work, we propose Text-to-Talk (TtT), a unified audio-text framework that integrates autoregressive (AR) text generation with non-autoregressive (NAR) audio diffusion in a single Transformer. By leveraging the any-order autoregressive property of absorbing discrete diffusion, our approach provides a unified training objective for text and audio. To support this hybrid generation paradigm, we design a modality-aware attention mechanism that enforces causal decoding for text while allowing bidirectional modeling within audio spans, and further introduce three training strategies that reduce train-test discrepancies. During inference, TtT employs block-wise diffusion to synthesize audio in parallel while flexibly handling variable-length outputs. Extensive experiments across Audio-QA and ASR tasks demonstrate the effectiveness of our approach, with detailed ablation studies validating each proposed component. We will open-source our models, data and code to facilitate future research in this direction.
@@ 1 Introduction
The recent success of LLMs has catalyzed a paradigm shift towards general-purpose Multimodal Large Language Models (MLLMs) capable of processing and generating information across diverse modalities (Xu et al., 2025; Team et al., 2023). Among these, speech-in/speech-out conversational systems have emerged as a pivotal component in facilitating natural human-AI interaction. Conventional systems typically decompose this problem into a cascaded pipeline of automatic speech recognition (ASR), LLM-driven response generation, and text-to-speech (TTS) synthesis. While effective to a degree, this modular design introduces significant latency accumulation and error propagation between modules, hindering naturalness and real-world applicability. In response, recent end-to-end approaches like Moshi Défossez et al. (2024), GLM4-Voice Zeng et al. (2024), and VITA-Audio Long et al. (2025) have sought to unify speech understanding and generation within a single model. These models are typically trained through multi-stage pipelines that involve text-to-audio tokenizer training, interleaved data construction, text-audio alignment and task-oriented supervised fine-tuning (Huang et al., 2025; Li et al., 2025; Ding et al., 2025; Chu et al., 2024). As shown in Fig. 1, these methods aim to generate interleaved text and speech tokens in an autoregressive manner, which are then decoded into continuous audio waveforms by a separate neural codec or diffusion-based decoder (Mehta et al., 2024; Kong et al., 2020).
Figure 1: (a) Distinct dependency structures for text and audio modality. (b) Due to disparate tokenization rates, the last audio span is of variable length.
@@
However, this emerging paradigm faces a fundamental challenges. As illustrated in Figure 1, we identify a fundamental mismatch in prevailing approaches that employ a single language model to autoregressively generate both text and audio tokens. This uniform treatment applies identical autoregressive training objectives across both modalities, overlooks a critical distinction in their underlying generative processes (Zeng et al., 2024; Xie & Wu, 2024; Borsos et al., 2023; Dang et al., 2024; Rubenstein et al., 2023). Text generation inherently follows a sequential causal structure characterized by strong target-target dependencies Box et al. (2015), where each token explicitly conditions on previously generated tokens. Consequently, an incorrect token prediction can propagate and introduce subsequent errors due to the exposure bias inherent in AR models Ranzato et al. (2015). In contrast, audio token generation is predominantly driven by source-target dependencies Ren et al. (2020), where audio output primarily condition on the source text rather than on the preceding audio tokens. Specifically, within the current NAR span, audio tokens generation should remain faithful to the source text even when previous audio tokens are incorrectly predicted. Applying a purely AR objective to audio generation thus introduces unnecessary sequential constraints, leading to suboptimal training dynamics and magnifying error propagation. This problem can be substantially alleviated by adopting a non-autoregressive generation strategy, which aligns better with the source-dependent nature of audio modeling. Recently, discrete diffusion has emerged as a compelling alternative to AR for discrete sequence modeling (Yu et al., 2025; Gong et al., 2024; Austin et al., 2021; Sahoo et al., 2024). Beyond empirical gains, recent theory shows that absorbing discrete diffusion can be interpreted as modeling the conditional distributions of clean tokens and admits a tight connection to any‑order AR objectives Ou et al. (2024).
Thus, we introduce Text‑to‑Talk (TtT), a unified audio‑text MLLM that integrates AR text generation with NAR audio diffusion within a single Transformer initialized from a pretrained LLM. Text segments are trained with a standard AR cross-entropy objective, while audio segments are modeled via an NAR discrete diffusion process. During inference, the model dynamically switches between AR and NAR decoding strategies based on special control tokens. In summary, our work makes the following contributions:
@@
  • We propose TtT, a hybrid AR-NAR MLLM that seamlessly integrates autoregressive text generation with discrete diffusion-based audio synthesis within a single Transformer initialized from a pretrained LLM. Our design preserves the reasoning and instruction-following capabilities of the base LLM while enabling efficient parallel audio generation.
@@ 2 Preliminary and notation
In this section, we establish the basic notation for interleaved audio-text sequences and provide brief overviews of the two core generative paradigms employed in our framework: AR modeling and absorbing discrete diffusion. These form the theoretical foundation of our proposed method in Section 3.
@@ 3.3 Multimodal Factorization and a Unified Objective
Intuitively, this averages the likelihood over all possible within-span orderings, reflecting the fact that audio tokens can be generated in any order given the appropriate cross-modal context. Note that while tokens within $\mathcal{A}_{m}$ form an antichain under the partial order (i.e., no mandatory sequential constraints), the order-marginalized conditional in Eq. 13 allows the model to leverage local target-target dependencies that may arise under specific generation orders. This flexibility enables the model to capture useful intra-span structures when beneficial.
@@
This formulation reveals that both modalities are fundamentally autoregressive: text uses a single linear extension (left-to-right), while audio integrates over all linear extensions consistent with the partial order.
@@
The key theoretical insight is that our combined training objective provides a tight upper bound on the negative log-likelihood of the desired joint distribution. To see this, consider the audio term:
@@
As illustrated in Figure 2 (a), our training pipeline starts from a pretrained text LLM and expands its vocabulary with discrete audio codebook tokens and control symbols ($\langle\text{SOA}\rangle$, $\langle\text{EOA}\rangle$). Each training sequence is organized as interleaved text spans and audio spans, where text spans are optimized with AR Loss while audio spans use the NAR diffusion objective.
Despite its theoretical and practical advantages, the hybrid AR-NAR paradigm introduces a significant train-test discrepancies that can degrade generation quality. During training, audio spans are partially masked according to the diffusion process, while during inference, the model must generate audio and text tokens conditioned on complete text context and previously generated clean audio tokens. To bridge this gap, we propose three principled training strategies:
@@ 4.2 Main Results
Table 1 main results part reports the comparison of conversational understanding ability across different models on Audio-QA task. From these results we make the following observations: (1) Our proposed TtT consistently achieves the best performance on all four Audio-QA datasets. For example, at the 3B scale, TtT attains scores of 17.46, 34.68, 6.53, and 11.61 on AlpacaEval, LLaMAQuestions, TriviaQA, and WebQuestions respectively, surpassing the Qwen2.5-Base (AR) model by +3.04, +24.68, +5.93, and +10.91. This improvement stems from the hybrid AR–NAR design, which better captures the inherent asymmetry between text and audio dependencies. (2) Our results demonstrate a clear scaling trend shared with the baselines, where larger backbones consistently yield stronger performance. For instance, TtT-3B achieves 34.68 on LLaMAQuestions, outperforming its 1.5B counterpart which scores 23.75, reflecting an absolute gain of +10.93. This pattern aligns with the scaling laws of LLMs, reflecting that additional capacity enhances both reasoning ability and cross-modal representation learning; (3) Qwen2.5-Base (NAR) performs noticeably worse than both our hybrid framework and the AR-only baseline. Since discrete diffusion models employ an order-agnostic AO-ARM objective that ignore the inherent sequential structure of interleaved text-audio sequences. This mismatch leads to order confusion, compromising sequence coherence and weakening audio-text alignment.